OpenStudy (aimmebgood):

factor completely x^(6) - y^(6). with explained steps please?

OpenStudy (jackellyn):

Try long division, take out (x-y) or (x+y)

OpenStudy (aimmebgood):

That doesnt help. If I knew how to solve this in any way, I would try it.

OpenStudy (jackellyn):

a^2 - b^2 = (a + b)(a - b) a^3 - b^3 = (a - b)(a^2 + ab + b^2) a^3 + b^3 = (a + b)(a^2 - ab + b^2) x^6 - y^6 = (x^3)^2 - (y^3)^2 = (x^3 + y^3)(x^3 - y^3) = (x + y)(x^2 - xy + y^2)(x - y)(x^2 + xy + y^2)